OpenStudy (anonymous):

i need 2 equations that have infintely many solutions

OpenStudy (anonymous):

@brandon125

OpenStudy (anonymous):

yea i can.

terenzreignz (terenzreignz):

Give me a linear equation. Involving x and y. Invent one, anything :)

OpenStudy (anonymous):

2x+4y

terenzreignz (terenzreignz):

equation... they have to be equal to something... any number, pick one :)

OpenStudy (anonymous):

=10

terenzreignz (terenzreignz):

Okay, 2x + 4y = 10 Now, pick a number, any number, except 1 or 0 :)

OpenStudy (nathan917):

2Y + 4X = 8 Y + 2X = 4

OpenStudy (nathan917):

*this is an example.*

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

okay, now multiply both sides of your equation 2x + 4y = 10 by 3

OpenStudy (anonymous):

6x+12=10

OpenStudy (anonymous):

*12y*

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

What about 10?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay, but I don't understand why...